OpenStudy (anonymous):

find the measure of an angle whose measure is 20 degrees more than its supplement

OpenStudy (anonymous):

supplementary angles are 180 degrees

OpenStudy (anonymous):

so its 180-x=20

OpenStudy (anonymous):

I would say 180 +20

OpenStudy (anonymous):

actually its 180-x=20

OpenStudy (tamtoan):

no teresaa, that angle x - 20 = 180 - x (it's supplement angle)

OpenStudy (anonymous):

maybe x+(x+20)=180

OpenStudy (anonymous):

The supplement of an angle takes it up to 180 degrees

OpenStudy (anonymous):

so how do i solve

OpenStudy (anonymous):

reading it literally... the measurement of an angle (x) is 20 more than its supplement x=180-(x+20)

OpenStudy (anonymous):

x=180-x-20

OpenStudy (anonymous):

2x=160

OpenStudy (anonymous):

You get to do the last step

OpenStudy (tamtoan):

eulergroupie, you are solving for x and x in your equation is ..."it's suplement", not the angle itself

OpenStudy (anonymous):

tamtoan can u tell me how u think its done

OpenStudy (anonymous):

It is strange wording...

OpenStudy (anonymous):

It is confusing... I would say that we have two angles that are supplementary. One is 100 degrees and one is 80 degrees. Together they make 180 degrees. That is the supplement. The angle that is 100 degrees is 20 more than 80 degrees. Sorry to be so confused... but I think that "the angle" that would be 20 more than its supplement would be 100 degrees. Whew!

OpenStudy (anonymous):

@tamtoan had a good point... my equation found the other angle.

OpenStudy (tamtoan):

teresass, follow eulergroupie and you will understand, he explained clearly there :)